40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog, Page 1 of 6  Not logged in ELOG logo
ID Date Author Type Category Subject
  257   Tue Feb 28 09:44:43 2023 AidanLab InfrastructureGeneralAdded lighting and tidied West Desk

I added LED stip lighting above the desk on the West wall. I also tidied that desk up (mostly in the drawers).

 

Attachment 1: IMG_3670.jpg
IMG_3670.jpg
  256   Thu Jan 12 11:52:47 2023 KojiLab InfrastructureGeneralHow to move the large engine hoist through the narrow door

How to move the large engine hoist through the narrow door

See http://nodus.ligo.caltech.edu:8080/Mariner/122

  255   Wed Jul 13 15:19:49 2022 JCElectronicsGeneralDesktop Computer

[JC]

The desktop computer is now running Debian Linux

  254   Fri Jul 8 07:06:41 2022 JcMiscVACDewar

[Stephen, JC]

Stephen and I added a new valve to Dewar's vacuum system. This valve allows the flow of atmospheric pressure into the system. We added 3 components to the system which were the valve, an adapter, and a T-intersection. After added these components, we continued to pump down only the highlighted yellow and green area with Dewar being closed off. The system pumped down to .1 mTorr until we decided to close off the pump. Once we close the pump off, we noticed the pressure began to rise. We took apart the system again and looked over the O-rings. We came across one ring with a sticky and clear material surround the rim and another ring with a fiber on it. We proceeded to clean and reassemble the system, but ran into this same issue. We tried to find where this leak was by squirting isopropyl around each ring and posibbly so a slower rise in pressure, but had no luck.

Afterwards, I checked both T-intersections individually and came across the same rise in pressure after closing off the valve for both tests. We suspect this may have been normal beforehand?

We continued by putting it all back together and taking data of the pumpdown.

Attachment 1: 29FCB85E-0AC8-43ED-8A43-31DC04459CD4_1_201_a.jpeg
29FCB85E-0AC8-43ED-8A43-31DC04459CD4_1_201_a.jpeg
  253   Thu Jun 23 10:53:17 2022 JCLab InfrastructureGeneralSeat For Photodiode Testing

[JC, Chub, Radhika]

Chub and I ordered a few parts from McMaster in order build a handrail-like stopper to keep the dewar from falling over. We also cut off the excess 8020 which was leaning over the table to fit. To hold down the support for the Dewar, Radhika and I decided to use C-clamps from the EE shop.

Attachment 1: 62AE8A4B-F817-4A03-A2F1-AE48B97FB935.jpg
62AE8A4B-F817-4A03-A2F1-AE48B97FB935.jpg
Attachment 2: IMG_1392_2.jpg
IMG_1392_2.jpg
  252   Fri May 20 16:42:47 2022 StephenThings to BuyOrderingUp to Air valve for IR Labs Cryostat

Purchased from Lesker, for venting of IR Labs Cryostat:

 - P/N C35103000 (KF10 Valve for up-to-air)

 - P/N QF25XQF10 (KF25-KF10 Reducer)

 - P/N QF25-100-T (KF25 Tee)

(to be connected in between the Chamber Tee and the Gate Valve)

Should be in next week!

  251   Wed May 11 19:46:21 2022 Aidan, Radhika, JordanLab InfrastructureVacuum chamberLeak cleaning on IR labs vacuum chamber - suspects

[Aidan, Jordan, Radhika]

Radhika and Jordan identified some particulates (hair and flecks of foil) on the O-ring on the IR labs dewer. Additionally, we saw a scratch in the O-ring groove and a nick on the metal of the base of the dewer where it meets the O-ring. All were in the leaky vicinity previously identified by the He testing.

We set up a cradle to hold the dewer while we are working on it. Still needs vertical supports.

R & J replaced the O-ring with a new one with Crytox applied. 

Attachment 1: IMG_8812.jpg
IMG_8812.jpg
Attachment 2: IMG_8799.jpg
IMG_8799.jpg
Attachment 3: IMG_8789.jpg
IMG_8789.jpg
Attachment 4: IMG_8805.jpg
IMG_8805.jpg
  250   Wed Apr 27 11:41:53 2022 AidanLab InfrastructureGeneralLab clean-up to get ready for PD testing in old TCS Lab

We (Aidan, Koji, Radhika, Aaron) partially tidied up the TCS Lab. The front table is clean and ready to recieve PD tesitng optics, electronics and vacuum hardware. We moved all electronics units (oscilloscopes, power supplies, etc) to the rack in the NE corner of the lab. The back table was partially tidied up. We need to schedule cleaning of the remaining tables in the lab and also an inventory and disposal of obsolete equipment in all the cupboards.

 

Attachment 1: IMG_8599.jpg
IMG_8599.jpg
Attachment 2: IMG_8598.jpg
IMG_8598.jpg
Attachment 3: IMG_8596.jpg
IMG_8596.jpg
Attachment 4: IMG_8595.jpg
IMG_8595.jpg
Attachment 5: IMG_8600.jpg
IMG_8600.jpg
  249   Thu Aug 26 09:00:39 2021 AidanMiscFloodLab flooded from broken pipe leading to sump room

Some photos of water and clean-up.

Summary: I came into the lab around 11:30AM and found water on the floor in the changing room outside QIL/TCS. Turns out the condesation overflow pipe from the AC blew out again. This time near the ceiling. Water was on the floor but also had sprayed a little onto the tool chest and East optical table. A few optics got wet on the table. Initial inspection looks like electronics were spared with the exception of the "broken" spectrum analyzer that was on the floor. 

Facilities came in and cleaned up the water. A small amount got into QIL but stayed near the door as the lab floor slopes up from the door area. They fixed the pipe and were looking into whether there was a blockage cuasing this problem. PMA was notified and John Denhart is coordinating follow-up.

Triage effort: given the AC was still active, John and I strung a temporary tarp across the two tables to block any spray.

Quote:

11:29AM - Lab has flooded again this morning. I'm calling PMA. Looks to be the same issue as before.

Quote:

 

Koji: QIL/TCS entrance flooding. Check your lab

Anchal: Can someone take a look at CTN too?

Koij: TCS needs more people @aidan

Koji: CTN ok

Aidan: On my way

Shruti: Cryo seems fine

Aidan: There was a leak in a pipe in the wall of B265A. It was coming from the building air conditioner condensation overflow. Facilities has fixed the pipe and is working on clean-up

 

 

Attachment 1: 20210825_-_TCS_Lab_Flood_-_compressed.pdf
20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf 20210825_-_TCS_Lab_Flood_-_compressed.pdf
  248   Wed Aug 25 11:28:56 2021 AidanMiscFloodLab flooded from broken pipe leading to sump room

11:29AM - Lab has flooded again this morning. I'm calling PMA. Looks to be the same issue as before.

Quote:

 

Koji: QIL/TCS entrance flooding. Check your lab

Anchal: Can someone take a look at CTN too?

Koij: TCS needs more people @aidan

Koji: CTN ok

Aidan: On my way

Shruti: Cryo seems fine

Aidan: There was a leak in a pipe in the wall of B265A. It was coming from the building air conditioner condensation overflow. Facilities has fixed the pipe and is working on clean-up

 

  247   Thu Aug 12 16:13:44 2021 AidanLab InfrastructureFloodCarpentry shop removed wet plaster sections from the wall - it's drying for a few days

The carpentry shop removed wet plaster sections from the wall following the flood (process was gentle scraping of wet plaster flakes, supervised by me). The wet section of wall needs a few days to dry and then they will plaster and paint it.

 

Attachment 1: IMG_4139.jpg
IMG_4139.jpg
  246   Tue Jul 27 08:39:33 2021 AidanMiscFloodTuesday (27-Jul) morning check - lab looks okay

I checked the lab this morning. It was dry and there wall was in the same state as yesterday.

Attachment 1: Image_Pasted_at_2021-7-27_08-38-2.jpeg
Image_Pasted_at_2021-7-27_08-38-2.jpeg
Attachment 2: Image_Pasted_at_2021-7-27_08-38.jpeg
Image_Pasted_at_2021-7-27_08-38.jpeg
  245   Mon Jul 26 16:23:03 2021 AidanMiscFloodLab flooded from broken pipe leading to sump room

 

Koji: QIL/TCS entrance flooding. Check your lab

Anchal: Can someone take a look at CTN too?

Koij: TCS needs more people @aidan

Koji: CTN ok

Aidan: On my way

Shruti: Cryo seems fine

Aidan: There was a leak in a pipe in the wall of B265A. It was coming from the building air conditioner condensation overflow. Facilities has fixed the pipe and is working on clean-up

Attachment 1: IMG_3881.MOV
  244   Tue Jun 15 08:47:27 2021 AidanLab InfrastructureGeneralCleaned up HWS table in preparation for lab move

I cleaned up the HWS table in preparation for replacement with the 4x10 table. We still need to move the cabinet and get the enclosure out of the way.

 

 

Attachment 1: IMG_3056.jpg
IMG_3056.jpg
  243   Wed Oct 30 12:13:14 2019 JonMiscDelivery NoteNew lab hardware

Ian and I moved some new hardware into the lab, shown in the below photos. It is from the shipment of loaned equipment recently returned by Whitman College. 

Attachment 1:

  • ZnSe 1" dia. lenses (x2)
  • ZnSe 38.1 mm dia. windows (x2)

Attachment 2:

  • Synrad 20 W CO2 laser
  • Synrad laser controller

Attachment 3:

  • Intraaction Model GE modulator driver
  • Intraaction AOM

Attachment 4:

  • Second Intraaction Model GE modulator driver
  • Unknown optical cavity (was packaged together with the driver)

The ZnSe lenses and windows were put in the CO2 drawer of the optics cabinet. The CO2 laser, AOM, and modulator drivers were left packaged in boxes underneath the large laser table.

Attachment 1: IMG_3782.jpg
IMG_3782.jpg
Attachment 2: IMG_3783.jpg
IMG_3783.jpg
Attachment 3: IMG_3784.jpg
IMG_3784.jpg
Attachment 4: IMG_3785.jpg
IMG_3785.jpg
  242   Tue Sep 3 10:49:57 2019 AidanLab InfrastructureGeneralExhaust duct added for new bake area in TCS lab

Facilities came in on Friday and teed off a new duct to provide exhaust for the proposed new vacuum bake area in the TCS Lab. Photos are attached. 

We installed a plastic sheet between the work area and the rest of the lab (the rest of the lab was overpressurized relative to the work area). Also, they use a vacuum when doing any drilling.

 

Attachment 1: IMG_1627.jpg
IMG_1627.jpg
Attachment 2: IMG_1629.jpg
IMG_1629.jpg
Attachment 3: IMG_1628.jpg
IMG_1628.jpg
Attachment 4: IMG_1626.jpg
IMG_1626.jpg
  241   Fri Aug 16 17:05:14 2019 Edita BytyqiElectronics FLIR Images of new reflector focusing heat

We got 11 new semi-circle cut reflectors of radius ~3.6 cm. I glued a screw to the back of one reflector using the same epoxy as for the previous reflectors. Due to the bigger ROC of the reflector, a tight focus is achievable at greater distances (~15 cm).

  240   Mon Aug 12 21:15:12 2019 Edita BytyqiElectronics Determining heater/reflector focus

I took images of the heat pattern projected on a piece of paper produced by the semi-circle reflector. I used 108V to drive current throught he heater. I tested the reflector without any coating and then with the dull and shiny sides of Al foil. I wasn't able to test the focal-point cut reflector because I had to glue a screw to it with epoxy which cures overnight. I will do these measurements tomorrow. Figure 2 shows the setup I used to get the data. The shiny side of Al foil is better at IR, so we will use that for the wavefront measurements.

Attachment 1: 20190812_151316.jpg
20190812_151316.jpg
Attachment 2: 20190812_151258.jpg
20190812_151258.jpg
  239   Fri Aug 9 20:33:11 2019 JonElectronicsRing HeaterPower-regulated heater ready to use

Despite some considerable time spent, I was not able to get the Omega SCR controllers working. The first unit definitely arrived damaged. None of its LED indicator lights ever functioned, despite those on the second controller working fine under the same setup. I tried swapping in the second controller, but it has no voltage output and a red LED is illuminated which, according to the manual, means "malfunction on trigger board" or "open SCR."  Either way, the remedy is to "consult factory."

Since we have get moving with data collection, I installed a simple variable transformer (borrowed from the 40m) which steps up/down the AC voltage from 0-120 V with the turn of a knob. I soldered the leads of one of the heaters to a standard power cable which plugs directly into the transformer. I have tested it and confirmed it to work.

Attachment 1: IMG_3576.jpg
IMG_3576.jpg
Attachment 2: IMG_3577.jpg
IMG_3577.jpg
  238   Tue Aug 6 19:22:28 2019 JonElectronicsComputingQIL NFS server set up

I've started building the NFS server for the QIL cymac. It's sitting on the workbench in the TCS lab next to the rack. Please don't move it or any of the parts behind it.

  237   Thu Aug 1 15:20:39 2019 Edita BytyqiLab InfrastructureOpticsReflector Mount and DC Supply

Yesterday, we were able to take some data using the 120 V DC power supply. The reflectors cut at the focal point and radius were both tested; the semi-circle cut proved to give a better focus, likely because roughly half the heat is lost using the focal-point reflectors. For upcoming tests, the semicircle reflectors will be used. We varied the surface shine by using the dull and reflective side of Al foil, as well as using the machined Al itself. The best result was given by using the more reflective side of Al foil.

Figure 1 shows the steady-state surface deformation profile detected by the HWS. The heaters don't have a uniform distribution along the wire, so more heat is radiated in the center of it, thus more of it is being focused to the center of the test optic. The data needs to be analyzed to determine the radius of the focus. Our rough estimate is about ~1.5 - 2 cm. We cannot collect any more data until we get a new power supply (AC 120 V).

Today, I came up with a new design for mounting the reflectors. I used a big 3-axis stage and a small 4-axis stage. This provides 5 degrees of freedom: 3 translational and 2 rotational, which is what we need for fine-tuning the focus and directing it at different angles incident to the test optic. The only problem with this design is that the 3-axis stage is too tall for the box, so the lid won't close.There is a smaller one available, but I have to figure out a way to increase its height, since the screw size is different from the ones on the pedestals available.

Additionally, Chub used metal-to-metal epoxy to glue a screw to the back of a reflector. I will wait until tomorrow to test it, because it is a slow acting epoxy. If it works, I have the necessary tools to do the same with the other reflectors. With the current deisgn the reflector wil be screwed in to where the round screw is in the stage. If it heats up a lot and affects the material of the stages, a small optical post (top of stage) will be used to make up for the absorbed heat.

 

Attachment 1: 20190731_170920.jpg
20190731_170920.jpg
Attachment 2: 20190801_145836.xcf
Attachment 3: 20190801_145836.jpg
20190801_145836.jpg
Attachment 4: 20190801_150635.jpg
20190801_150635.jpg
Attachment 5: 20190801_145842.jpg
20190801_145842.jpg
  236   Mon Jul 29 18:53:16 2019 Edita BytyqiLab InfrastructureOpticsMounted Reflector and Heater

Since we set up the 2-lens system focusing the laser beam to the CCD, the next step was to mount the spherical reflector (31 mm wide) and the heater (~3 mm diameter). I used a small 3-axis stage to mount the heater, providing 3 degrees of freedom that would allow to manipulate the height of the heater, its position with respect to the reflector (left-right and in-out). The reflector was mounted in such a way that we can control its rotation angle, height and horizontal displacement. The current design is not quite sophisticated as it is just a first test, however I will look into different tools in the lab to see if I can use less mounts to get the same degrees of freedom.

The new heaters are supposed to be heated using AC. We used a DC power supply and ran ~30V through the wire, however only about ~50 mA of current was running through it. Jon will look into the specs of the new heaters to see if the power supply was the problem.

Attachment 1: 20190729_185214.jpg
20190729_185214.jpg
Attachment 2: 20190729_185156.jpg
20190729_185156.jpg
Attachment 3: 20190729_185153.jpg
20190729_185153.jpg
Attachment 4: 20190729_185139.jpg
20190729_185139.jpg
Attachment 5: 20190729_185112.jpg
20190729_185112.jpg
  235   Thu Jul 25 09:13:36 2019 JonThings to BuyGeneralNeed to Order Gloves

New gloves are ordered for the TCS and QIL labs. They arrive tomorrow (Friday).

Quote:

Small/Medium size gloves need to be ordered in order to handle the optics carefully.

 

  234   Wed Jul 24 16:25:13 2019 Edita BytyqiLab InfrastructureOpticsUpdated 2-lens setup

The previous 2-lens setup focused the beam to a tight spot, however due to the divergence angle of the laser beam, a significant amount of power was not being captured by the fiirst lens at a distance of 40 cm from the source. The divergence angle seems to be bigger than 0.06 by a factor of 2, so a f = 20 cm lens was used to collimate the beam and a f = 30 cm lens was used to focus it. A mirror was used to reflect the beam, so we obtain steering control. Additionally, the focusing lens was placed on a small 1-axis stage in order to control the distance of the lens from the CCD, providing control over the focused beam size.

Note: The 30 cm lens was cleaned with methanol, however it still has some residue on the surface. The beam imaged to the Harrtmann Sensor looks good, however the lens will be cleaned by using a different solvent or replaced by a different 30 cm lens. The 3 lenses at the edge of the box will stay inside in order to prevent contamination, however they will not be used in the design.

Attachment 1: 20190724_161926.jpg
20190724_161926.jpg
  233   Mon Jul 22 18:46:23 2019 Edita BytyqiLab Infrastructure Laser-Lens-HWS Setup

Today, I set up a system consisting of the 520 nm laser, a 2'' mirror and two lenses of focal lengths f1 = 40 cm and f2 = 20 cm. The goal was to collimate the beam coming from the laser, so it goes parallel through the test optic at a radius of ~2.5 cm and then focus it to a radius of ~ 1.2 cm to fit the CCD dimensions of the HWS. The mirror was placed about 1 cm close to the laser and the first lens is setup at a distance~f1=40cm from the mirror. The test optic is placed between the two lenses and the second lens is placed about 10 cm from the CCD. The distance between the two lenses isn't important and could change in the future. The lenses and mirrors are all labeled.

I measured the approximate angle of divergence (0.06 rad) of the laser by taking the beam diameter at different positions along the propagation axis. This allowed for the ABCD matrix calculations to be finalized and the focal lengths of the lenses be chosen accordingly. 

In order to have more space in the box, I removed everything that was not necessary to the side.

Attachment 1: 20190722_183933.jpg
20190722_183933.jpg
Attachment 2: 20190722_183928.jpg
20190722_183928.jpg
  232   Mon Jul 22 18:44:53 2019 Edita BytyqiThings to BuyGeneralNeed to Order Gloves

Small/Medium size gloves need to be ordered in order to handle the optics carefully.

  231   Mon Jul 22 09:44:21 2019 JonLab InfrastructureComputingInstalled Amplifying VGA Splitter

The VGA signal outputted by the multiplexer is too weak to drive two monitors. This has required video cables to be manually switched back and forth between the monitor mounted above the laser table and the desktop console.

Today I solved this problem by installing an amplifying (active) VGA splitter on the video output of the multiplexer. One output of the amplifier goes to the desktop monitor and the other to the laser table. We can now monitor the HWS realtime GUIs directly above the optical setup, with no cable swapping.

Attachment 1: IMG_3522.jpg
IMG_3522.jpg
Attachment 2: IMG_3421.jpg
IMG_3421.jpg
  230   Tue Jun 18 19:05:16 2019 JonLab InfrastructureGeneralLarge optical table cleared and set up for point absorbers

Aidan and I continued the lab clean-up today. There's still more to do, but we did fully clear the large optical table which formerly housed the 50 W CO2 laser. I moved the optical enclosure over from the small table to serve as the area for the point absorber experiment. Inside it I mounted the Hartmann sensor and a 532 nm Thorlabs LED source. The LED still needs collimating/focusing optics to be installed.

Attachment 1: IMG_3421.jpg
IMG_3421.jpg
  229   Mon Jun 17 20:29:34 2019 JonLab InfrastructureComputingServer rack cleaned up, new workstation installed

In anticipation of the point absorber SURF project, I cleaned up the server rack and installed a new workstation today.

The workstation replaces the old one, whose hard drive had failed, with a more powerful machine. The hostname (tcs-ws), IP address (10.0.1.168), user name (controls), and standard password (written in the secret place) are all the same as before.

I moved the control consol from its old spot in the back corner of the lab to the bench beside the rack. This is a more convenient location because the Hartmann sensor realtime GUIs can now be easily seen from the optical tables. I mounted the HWS machine in the rack as well and reconnected the video multiplexer to all the machines.

I tested the Hartmann sensor Python software and confirmed it to be working. It required a minor bug fix to the realtime gradient field GUI code. It seems that since this script was last run, the input data file type has switched from pickled numpy to HDF5.

Attachment 1: IMG_3419.jpg
IMG_3419.jpg
  228   Thu May 23 11:36:49 2019 aguptaLab InfrastructureBorrowedBorrowed Label Maker

I have borrowed TCS's label maker in CTN for few days. If you need it, you can take it from the top of blue cabinets.

  227   Mon Apr 15 08:12:58 2019 aaronLab InfrastructureGeneralMeasurements for Bake Station

My preference is to have tubes back towards the wall where possible. We might be able to drill a large diameter hole in the table top to accommodate them.

We have to get confirmation that the exhaust can be extracted - otherwise this whole thing is moot.

Quote:

I drew up one way we could set up the three available bake ovens in the TCS lab on the single oil pump. 

If this looks feasible to others, we can move the ovens into the TCS lab. Duo and I will be occupied at KNI during Tuesday and Monday morning, so the usual lab cleanup time may not be the best. Perhaps Monday afternooon we can at least get the ovens out of the hallway and get one of them set up for baking.

 

  226   Fri Apr 12 17:40:58 2019 aaronLab InfrastructureGeneralMeasurements for Bake Station

I drew up one way we could set up the three available bake ovens in the TCS lab on the single oil pump. 

If this looks feasible to others, we can move the ovens into the TCS lab. Duo and I will be occupied at KNI during Tuesday and Monday morning, so the usual lab cleanup time may not be the best. Perhaps Monday afternooon we can at least get the ovens out of the hallway and get one of them set up for baking.

Attachment 1: BakeStation_Page_1.jpg
BakeStation_Page_1.jpg
Attachment 2: BakeStation_Page_2.jpg
BakeStation_Page_2.jpg
  225   Fri Feb 8 10:48:33 2019 AidanLab InfrastructureGeneralWater damage repair work in the TCS Lab

Caltech Facilities has determined that the walls in the SE corner of the TCS Lab in West Bridge were water damaged during last weekend’s rain. They are going to remove the plaster from the walls and dehumidify the area for a week or so. All tables in the room are going to be covered with plastic for this process. In the short term I’ve shutdown all the equipment in the lab (including FB4). The 2-micron cavity-testing fabrication has been moved next door to the QIL.

  224   Tue Dec 4 16:57:08 2018 AidanComputingHartmann sensorUpdated GIT version of HWS code

I changed the HWS code to the new git.ligo HWS version.

  • Object files are in ~/hws
  • scripts are in ~/hws-server
  • utilities are still in old git repo moved to ~/.HWS_code_temporary_home

I've set up some symbolic links to these directories to mimic the old directory structure, so ..

  • ~/pyHWS links to new object file directory
  • ~/pyHWS/scripts lnks to new scripts directory
  • ~/pyHWS/utilities links to old utilities directory
  223   Fri Oct 26 09:49:47 2018 AidanMiscOpticsStarting 80C cure of epoxy

I've started an 80C cure of two materials bonded by EPOTEK 353ND. The objective is to see (after curing) how much the apparent glass transition temperature is increased over a room-temperature cure.

Attachment 1: IMG_7054.jpg
IMG_7054.jpg
  222   Wed Oct 17 16:17:45 2018 awade, AidanComputingDAQAdded AWC thermocouple reader back to acromag1

Acromag IOC process was removed from PSL lab acromag1 computer a few months ago.  Aidan needs them again but it would be better if it were run from TCS lab computers.

An instance of the modbus IOC is now running on tcs-ws within a docker container.  Docker is named tcslabioc.  Configuration files are located in ~/modbus.  Instructions on how to use the docker are located in ATF:2249.  To install docker see google.

To set up the specific instance in the TCS lab run

>sudo docker run -dt --name tcslabioc -v /home/controls/modbus/test_acromag.cmd:/home/modbus/IOCStart.cmd -v /home/controls/modbus:/home/modbus -p 5064:5064 -p 5065:5065 -p 5064:5064/udp -p 5065:5065/udp andrewwade/modbusepicsdocker 

Then whenever you want to stop, run:

> sudo docker stop tcslabioc

or to restart run

>sudo docker restart tcslabioc. 

So if you update the .cmd or .db files just run the restart command above and the channels should automatically update when it reboots. For other cleanup and control commands see docker documentation.  It can also be configured to keep alive on system reboot.

The cmd and db files are included below in the attachments for reference. 

 

Quote:
  • Thermcouple reader is plugged into an Acromag ADC (10.0.1.55)
  • acromag1 is now reading this in:
    • ~/modbus/iocBoot/iocTest/test_acromag.cmd (contains ip address of ADC)
    • ~/modbus/db/TEMPMONCHANS.db (contains the EPICS channel definitions and calibration of thermocouple reader)
  • C4:AWC-TEMPMON_C (output channel)

Restored work done in http://nodus.ligo.caltech.edu:8080/TCS_Lab/201

 

 

 

 

Attachment 1: TCSIOCFILS.tar.gz
  221   Thu Aug 16 11:52:27 2018 AidanComputingDAQAdded AWC thermocouple reader back to acromag1
  • Thermcouple reader is plugged into an Acromag ADC (10.0.1.55)
  • acromag1 is now reading this in:
    • ~/modbus/iocBoot/iocTest/test_acromag.cmd (contains ip address of ADC)
    • ~/modbus/db/TEMPMONCHANS.db (contains the EPICS channel definitions and calibration of thermocouple reader)
  • C4:AWC-TEMPMON_C (output channel)

Restored work done in http://nodus.ligo.caltech.edu:8080/TCS_Lab/201

 

 

 

  220   Fri Aug 3 15:46:12 2018 Aria ChaderjianLaserGeneralAugust 3, 2018

To the best of my ability, calculated the magnification of the plane of the test optic relative to the HWS (2.3) and input this value.

Increased the temperature slightly and saved data points of defocus to txt files when temperature leveled out. This was a slow process, as it takes a while for things to level out. I only got up to about 28.5C, and will need to continue this process.

I also plotted the best-fit defocus for each temperature from COMSOL (Temperature vs. Defocus), and looking at values from HWS it seems that we're off by a normalization factor of approx. 4.

  219   Tue Jul 24 16:52:44 2018 Aria ChaderjianLaserGeneralJuly 23, 2018 and July 24, 2018

Attached the grid array of the HWS.

Applied voltage (5V, 7V, 9.9V, 14V) to the heater pad and took measurements of T and spherical power (aka defocus).

The adhesive of the temperature sensor isn't very sticky. The first time I did it it peeled off. (Second time partially peeled off). We want to put it on the side of Al if possible.

Bonded a mirror (thickness ~6 mm) to aluminum disk (thickness ~5 mm) and it's still curing.

  218   Mon Jul 23 10:04:19 2018 Aria ChaderjianLaserGeneralJuly 20, 2018

With Jon's help, I changed the setup to include a mode-matching telescope built from the f=60mm (1 inch diameter) lens and the f=100mm lens. These lenses are located after the last gold mirror and before the test optic. The height of the beam was also adjusted so that it is more centered on these lenses. Note: these two lenses cannot be much further apart from each other than they currently are, or the beam will be too large for the f=100mm lens.

We considered different possible mounts to use for the test optic, and decided to move it to a mount where there is less contact. The test optic was also moved closer to the HWS to achieve appropriate beamsize on the optic coming from the mode-matching telescope.

The f=200 lens is now approximately 2/3 of the distane from the test optic to the HWS, resulting in an appropriately sized beam at the HWS.

Current was also turned down to achieve 0 saturated pixels.

Attachment 1: 37767216_2077953269190954_5080703241788850176_n.jpg
37767216_2077953269190954_5080703241788850176_n.jpg
  217   Fri Jul 13 16:42:50 2018 Aria ChaderjianLaserGeneralJuly 13, 2018

The table is set up. The HWS and SLED were moved slightly, and a minimal angle between the test mirror and HWS was achieved.

There are two possible locations for the f=60mm lens that will achieve appropriate magnification onto the HWS: 64cm or 50 cm from the f=200mm lens. 

At 64cm away, approximately 79000 saturated pixels and 1054 average value.

At 50cm away, approximately 22010 saturated pixels and 1076 average value.

Currently the setup is at 64cm. Could afford to be more magnified, so might want to move the f=60mm lens around. Also, if we're going to need to be able to access the HWS (i.e. to screw on the array) we might want to move to the 50cm location.

  216   Thu Jul 12 18:48:21 2018 Aria ChaderjianLaserGeneralJuly 12, 2018

The beam reflecting off the test mirror was clipping the lens between gold mirror and test mirror, so I reconfigured some of the optics, unfortunately resulting in a larger angle of incidence.

From the test mirror, the beam size increases much too rapidly to fit onto the 2-inch diameter lens with f=8 that was meant to resize the beam for the CCD of the HWS. It seems that the f=8 lens can go about 6 inches from the test mirror, and an f ~ 2.3 (60 mm) lens can go about 2 inches in front of the CCD to give the appropriate beam size. However, the image doesn't seem very sharp.

The beam is also not hitting the CCD currently because of the increase in angle of incidence on the test mirror and limitations of the box. I'd like to move the HWS closer to the SLED (and will then have to move the SLED as well).

  215   Tue Jul 10 17:49:13 2018 Aria ChaderjianLaserGeneralJuly 10, 2018

Went down to the lab and showed Rana the setup. He's fine with me being down there as long as I let someone know. He also recommended using an adjustable mount  (three screws) for the test mirror instead of the mount with top bolt and two nubs on the bottom - he thinks the one with three screws as constraints for the silica will be easier to model (and be more symmetric constraints)

Mounted the f=8" lens (used a 2" pedestal) and placed it on the table so the image fit well on the CCD and so a sharp object in front of the lens resulted in a sharp image. The beam was clipping the f=4" lens (between gold mirror and test mirror) so I spent time moving that gold mirror and the f=4" lens around. I'll still need to finish up that setup.

 

Attachment 1: IMG_4518.jpg
IMG_4518.jpg
Attachment 2: IMG_4517.jpg
IMG_4517.jpg
  214   Thu Jun 21 17:12:38 2018 AidanLaserHartmann sensorTwo inch PBS from Edmund Optics - effect on ETM HWS transmission

I'm considering the 86-711 2" 532nm PBS from Edmund Optics for the ETM HWS at the sites.

https://www.edmundoptics.com/optics/polarization/linear-polarizers/532nm-50mm-diameter-high-energy-laser-line-polarizer/

The effect on the transmission through the system, compared to the THorlabs PBS, is shown in the attached plot.

Conclusion: it looks almost as effective as the Thorlabs PBS with the added benefit of being 2" in diameter.

Attachment 1: PBS_Thorlabs_v_EdmundOptics.pdf
PBS_Thorlabs_v_EdmundOptics.pdf
  213   Wed May 30 18:59:49 2018 awadeComputingDAQFB4 ip address has changed

Maybe you've resolved this now. I moved the DHCP allocation to 10.0.1.160 and above around that time because a bunch of misc devices were starting to populate the dynamically allocated IP space around there.  I'd say FB4 was not setup with a manual IP at the time.

 

Quote:

We've had trouble logging into FB4. I access the computer directly in the AWC lab and found that the IP address had changed from 10.0.1.156 to 10.0.1.161.

I'm not sure how this happened. It's possible that the IP address is not set to a static value and FB4 was rebooted. I'm not familar with Debian so I don't know where to look to find whether the IP address is static or not.

The DAQD is still running.

 

  212   Wed May 30 18:50:05 2018 awadeComputingDAQRebooted fb4

Time was still off by nine days as of yesterday.  I tried rebooting remotely to see if time would correct to system clock.  It didn't and fb4 hung.

Just manually restarted the box.  Now dataviewer is showing a 'Time Now'  of 5 Jan 1980.   

Not sure how to set the frame builder clock time. Ultimately the best solution is to have ADC cards but can we find a hack for now? Is it possible to run a cron script it to reset time to the computer time?

 

 

 

Quote:

The framebuilder on FB4 thinks the current time is 26-Jan-2018 6:18AM UTC. The date command on FB4 yields the correct date and time (5-Feb-2018 15:17 PST).

There is a major error with the framebuilder clock.

 

  211   Fri May 18 16:32:37 2018 AidanLaserHartmann sensorHWS green LED fiber launcher - D1800125-v5_SN01

Title was wrong - this is actually config [12,2,4,125]

Quote:

Here is the output from D1800125-v5_SN01.

Quote:

And here's the output of the fiber launcher when I fixed it at 313mm from the camera, attached an iris to the front and slowly reduced the aperture of the iris.

The titles reflect the calculated second moment of the intensity profiles (an estimate of the equivalent Gaussian beam radius). The iris is successful in spatially filtering the central annular mode at first and then the outer annular mode. 

We'll need to determine the optimum diameter to get good transmission spatially without sacrificing too much power.

Quote:

 [Aidan, Marie]

We tested the output of the fiber launcher D1800125-v3. We were using a 6mm spacer in the SM1 lens tube and 11mm spacer in the SM05 lens tube and the 50 micron core fiber.

The output of the fiber launcher was projected directly onto the CCD. Images of these are attached (coordinates are in pixels where 100 pixels = 1.2mm)

There is a lot of high-spatial frequency light on the output. It looks like there is core and cladding modes in addition to a more uniform background. There was an indication that we could clear up these annular modes with an iris immediately after the fiber launcher but I didn't get any images. We're going to test this next week when we get an SM1 mountable iris.

 

 

 

  210   Wed May 16 16:47:29 2018 AidanLaserHartmann sensorHWS green LED fiber launcher - D1800125-v5_SN01

Here is the output from D1800125-v5_SN01.

Quote:

And here's the output of the fiber launcher when I fixed it at 313mm from the camera, attached an iris to the front and slowly reduced the aperture of the iris.

The titles reflect the calculated second moment of the intensity profiles (an estimate of the equivalent Gaussian beam radius). The iris is successful in spatially filtering the central annular mode at first and then the outer annular mode. 

We'll need to determine the optimum diameter to get good transmission spatially without sacrificing too much power.

Quote:

 [Aidan, Marie]

We tested the output of the fiber launcher D1800125-v3. We were using a 6mm spacer in the SM1 lens tube and 11mm spacer in the SM05 lens tube and the 50 micron core fiber.

The output of the fiber launcher was projected directly onto the CCD. Images of these are attached (coordinates are in pixels where 100 pixels = 1.2mm)

There is a lot of high-spatial frequency light on the output. It looks like there is core and cladding modes in addition to a more uniform background. There was an indication that we could clear up these annular modes with an iris immediately after the fiber launcher but I didn't get any images. We're going to test this next week when we get an SM1 mountable iris.

 

 

Attachment 1: HWS_LED_D1800125_V5_sn01.pdf
HWS_LED_D1800125_V5_sn01.pdf
Attachment 2: HWS_LED_D1800125_V5_sn01.m
% get the beam size from the HWS ETM source D1800125-v5_sn01

[out,r] = system('tar -xf HWS*.tar');

% load the files
dist = [1,10,29,51,84,105,140,180,240,295,351,435,490,565]; % beam propagation distance
files = dir('*.raw');
close all

... 49 more lines ...
Attachment 3: HWS_D1800125-v5_sn01_characterize_20180516.tar
  209   Fri May 11 16:18:47 2018 AidanLaserHartmann sensorHWS green LED fiber launcher

And here's the output of the fiber launcher when I fixed it at 313mm from the camera, attached an iris to the front and slowly reduced the aperture of the iris.

The titles reflect the calculated second moment of the intensity profiles (an estimate of the equivalent Gaussian beam radius). The iris is successful in spatially filtering the central annular mode at first and then the outer annular mode. 

We'll need to determine the optimum diameter to get good transmission spatially without sacrificing too much power.

Quote:

 [Aidan, Marie]

We tested the output of the fiber launcher D1800125-v3. We were using a 6mm spacer in the SM1 lens tube and 11mm spacer in the SM05 lens tube and the 50 micron core fiber.

The output of the fiber launcher was projected directly onto the CCD. Images of these are attached (coordinates are in pixels where 100 pixels = 1.2mm)

There is a lot of high-spatial frequency light on the output. It looks like there is core and cladding modes in addition to a more uniform background. There was an indication that we could clear up these annular modes with an iris immediately after the fiber launcher but I didn't get any images. We're going to test this next week when we get an SM1 mountable iris.

 

Attachment 1: LED_test_02_20180511.pdf
LED_test_02_20180511.pdf
  208   Fri May 11 12:32:21 2018 AidanLaserHartmann sensorHWS green LED fiber launcher

 [Aidan, Marie]

We tested the output of the fiber launcher D1800125-v3. We were using a 6mm spacer in the SM1 lens tube and 11mm spacer in the SM05 lens tube and the 50 micron core fiber.

The output of the fiber launcher was projected directly onto the CCD. Images of these are attached (coordinates are in pixels where 100 pixels = 1.2mm)

There is a lot of high-spatial frequency light on the output. It looks like there is core and cladding modes in addition to a more uniform background. There was an indication that we could clear up these annular modes with an iris immediately after the fiber launcher but I didn't get any images. We're going to test this next week when we get an SM1 mountable iris.

Attachment 1: LED_test_01_20180511.pdf
LED_test_01_20180511.pdf
ELOG V3.1.3-